OpenStudy (anonymous):

Find the derrivative of (3-4e^x)/x^3 So far I have gotten ((-4e^x) (x^3)-3(3x^2))/x^6 is this correct?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

close but the second term in the numerator is wrong

OpenStudy (anonymous):

\[\left(\frac{f}{g}\right)'=\frac{f'g-g'f}{g^2}\]

OpenStudy (anonymous):

\(f(x)=3-4e^x,f'(x)=-4e^x,g(x)=x^3,g'(x)=3x^2\)

OpenStudy (anonymous):

unless you did some algebra maybe lets see

OpenStudy (anonymous):

\[\frac{x^3\times (-4e^x)-3x^2(3-4e^x)}{x^6}\] is the first step

OpenStudy (anonymous):

i was wrong, cancel a common factor of \(x^2\) and get \[\frac{x\times (-4e^x)-3(3-4e^x)}{x^4}\]

OpenStudy (anonymous):

then multiply out in the numerator

OpenStudy (anonymous):

Ohh..I think I tried taking out the -4e^x on the numerator..

OpenStudy (anonymous):

So would the answer be x(-4e^x)-9+12e^x?

OpenStudy (anonymous):

that looks beter

OpenStudy (anonymous):

for the numerator

OpenStudy (anonymous):

Alright thanks a lot you were a lot of help! :) Annd yess it'd be all over x^4 :) I almost forgot that part ha

OpenStudy (anonymous):

yw